OpenStudy (anonymous):

Find the slope of a line passing through (6, 7) and (2, 9) A. negative one-half B. one-fourth C. -2 D. -4

OpenStudy (hoblos):

A

OpenStudy (anonymous):

rise over run = y/x = -2/4 = -1/2 therefore answer = A